OpenStudy (anonymous):

For the following system, if you isolated y in the first equation to use the Substitution Method, what expression would you substitute into the second equation? 4x − y = 8 2x + 3y = 9 8 − 4x 8 + 4x −8 + 4x −8 − 4x

OpenStudy (anonymous):

\[4x-y=8\] \[4x=y+8\] \[4x-8=y\]

OpenStudy (anonymous):

since \(y=4x-8\) you would substitute \(4x-8\) for \(y\) in the second expression.

OpenStudy (anonymous):

solve \[2x + 3(4x-8) = 9\] for \(x\)

OpenStudy (anonymous):

6z -24 = 9?